OpenStudy (anonymous):

5808= 1/2*k*1.2^2 Solve for K

OpenStudy (anonymous):

is that 1/(2*k*1.2^2) or (1/2)*k*1.2^2

OpenStudy (anonymous):

k=5808*2/(1.2^2)=8.0667*10^3 assuming 5808= (1/2)*k*1.2^2

OpenStudy (anonymous):

or k = 1.2^2/(5808*2) if 5808= (1/2*k)*1.2^2

OpenStudy (anonymous):

and finally k = 1/(5808*2*1.2^2) if 5808= 1/(2*k*1.2^2)

OpenStudy (anonymous):

k=8066.666667

OpenStudy (mayankdevnani):